How many grams of NO will be produced, if 1.93g of H2O are also produced in the reaction below? Show the solution?




4NH3 + 5O2 -----> 4NO + 6H2O

1
Expert's answer
2022-06-01T20:17:03-0400

Mr (NO) = 30 g/mol;

Mr (H2O) = 18 g/mol.


4NH3 + 5O2 = 4NO + 6H2O

According to the above reaction, 6 mol (108 g) of H2O is formed when 4 mol (120 g) of NO is formed. We find how many grams of NO are formed when 1,93 g of H2O are formed:

120g NO — 108g H2O

X — 1,93g H2O

X = 120 * 1,93 / 108 = 2,144 g

ANSWER: 2,144 g NO